OpenStudy (anonymous):

cylinder a has a radius 1 m and height 4 m. cylinder b has radius 2 m and height 4 m. What is the ratio of volume cylinder a to cylinder b?

OpenStudy (anonymous):

1Cylinder A volume is 12.57 and cylinder b volume is 50.27 how do i make it into a ratio

ganeshie8 (ganeshie8):

Volume of cylinderA = \(\large \pi * 1^2*4 = 4\pi \) Volume of cylinderB = \(\large \pi *2^2*4 = 16\pi \)

ganeshie8 (ganeshie8):

To get the ratio of volumes of A to B, just divide : ratio of A to B = \(\large \dfrac{4\pi}{16\pi} = \dfrac{1}{4}\)

OpenStudy (anonymous):

THANK YOU!!!!!!!!!!!!!!!!!!!!!!!!!!!

ganeshie8 (ganeshie8):

u wlc :)

ganeshie8 (ganeshie8):

btw dividing 12.57 by 50.27 must also give u the same ratio 1/4